# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 6. Geometric data
$\int_{\mathbb{H}}a^{*}(z^{\prime})\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime}.$
The following lemma deals with the integration over the part of the fundamental domain with large injectivity radius, giving us the upper bound consisting of two terms: one is an integral with integrand with exponential decay and the other is an integration over the points with the small injectivity radius. The first term is desired and can be united with the bound acquired from the small injectivity radius case, but the second term needs to be further bounded. This is taken care of when we deal with all the parts dealing with the parts with small injectivity radius, which is done in Lemma 6.5.
@@
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. $D(4T)$ denotes the points with injectivity radius bigger than $4T$. Let $a$ be a mean zero function on $X$. Then
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right|$
$=\ \|a\|_{2}^{2}\,\int_{0}^{2T}\,11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}\,(1+\rho)\,e^{-\beta\rho}\,d\rho$
$+\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|,$ (6.7)
where $\beta=\beta(\lambda_{1}(X))$, $\lambda_{1}(X)$ is the spectral gap of Laplace-Beltrami operator $\Delta_{X}$, and where
The proof is based on modifying the expression so that the exponential mixing theorem, Theorem 1.2, can be applied.
@@
As $z\in D(4T)$, we have that $d(z,\gamma z)$ is larger than $4T$ for any non-identity group element $\gamma$. We also have that $z^{\prime}\in B(z,2T)$. This means that $d(\gamma z,z^{\prime})$ is larger than $2T$ for any non-identity group element. As $F_{t,t^{\prime},d(z,z^{\prime})}$ equals 0 if $d(z,z^{\prime})$ is larger than $2T$, we have that only the identity element can have a non-zero contribution. Expression (6.4) equals
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.8)
By adding and subtracting
$\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz$
in the absolute values of Expression (6.8) and then appyling the triangle inequality gives the following expression as an upper bound for Expression (6.8):
$\left|\int_{D}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.9)
Focus on the first term. We write $z^{\prime}$ using polar coordinates around $z$, so $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, and the first term equals
$\left|\int_{D}a(z)\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)\,a^{*}(\varphi_{\rho}(z,\theta))\,F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz\right|,$
where $a$ can be seen as a function on the unit tangent bundle by defining that $a(z,\theta)=a(z)$. The integrals of this expression can be reordered using Fubini’s theorem yielding the following expression:
$\left|\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\int_{D}\int_{\mathbb{S}^{1}}a(z)\,a^{*}(\varphi_{\rho}(z,\theta))\,d\theta\,dz\,d\rho\right|.$
By pulling the absolute values inside the first integral, we have the following upper bound, where the inner two integrals are written as an inner product:
@@
We can apply the Exponential mixing theorem, Theorem 1.2, to the absolute value of the inner product which then yields the following upper bound for the previous expression:
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}(1+\rho)e^{-\beta\rho}\,d\rho.$
Hence Expression (6.9) and hence the original expression we set out to bound are bound by
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,e^{\beta}\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}d\rho\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$
The following lemma deals with the integration over the part of the fundamental domain with small injectivity radius giving an upper bound for it. This can then be united with the bound acquired for the large injectivity radius case in Lemma 6.4.
@@
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. $D(4T)$ denotes the points with injectivity radius larger than $4T$. Then
$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
is bounded from above by
$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho$
The proof is based on careful use of Hölder’s inequality together with a careful management of the sum elements with non-zero contribtution.
@@
$\left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a^{*}(z^{\prime})F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|.$
By pulling the absolute values inside this is bounded by
$\int_{D\setminus D(4T)}|a(z)|\int_{B(z,2T)}\left|a^{*}(z^{\prime})\right|F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz,$
which is further bounded by the following expression when using the supremum norm bounds for $|a|$ and $|a^{*}|$:
$\|a\|_{\infty}^{2}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$
In the previous expression, the sum elements with a possible non-zero contribution are those $\gamma\in\Gamma$ such that $d(\gamma z,z^{\prime})$ is smaller than $2T$ for some $z^{\prime}\in B(z,2T)$. This collection depends on $z$ and we denote it by $\Gamma(z)$. We have that
$\#\Gamma(z)\leq\frac{|B(z,4T)|}{\operatorname{InjRad}_{X}}\leq 2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}},$
which follows from that the minimum distance of $\gamma z$ and $\gamma^{\prime}z$, where $\gamma\neq\gamma^{\prime}$ is $\operatorname{InjRad}_{X}$, and that $\gamma z$ must lie in $B(z,4T)$ for it to be possible that $d(\gamma z,z^{\prime})$ is smaller than $2T$, since $z^{\prime}\in B(z,2T)$. Hence the previous integral expression has the following upper bound when using Hölder’s inequality and $\Gamma(z)$ instead of $\Gamma$:
$\|a\|_{\infty}^{2}\,\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}\right)^{\frac{1}{2}}\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}\left(\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}\right)^{2}\right)^{\frac{1}{2}}.$ (6.10)
Focus on the double integral in the second factor. Using Hölder’s inequality to sum and using the bound observed for $\Gamma(z)$ gives the following upper bound for it:
@@
$2\pi\,\int_{D\setminus D(4T)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz.$
By using that $d(\gamma z,z^{\prime})$ equals $d(z,\gamma^{-1}z^{\prime})$ and interchanging the order of the sum and the integral using the linearity of integral, the previous expression equals
@@
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,\gamma^{-1}z^{\prime})}^{2}dz^{\prime}dz,$
which further equals under a change of variables $z^{\prime}\mapsto\gamma z^{\prime}$ the following:
@@
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
Due to the presence of $F_{t,t^{\prime},d(z,z^{\prime})}$, for non-zero contribution, we must have that $d(z,z^{\prime})$ is less than $2T$. This means that the previous expression equals
@@
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)\cap B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz,$
which is then bounded by
$2\pi\,\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}\frac{e^{4T}}{\operatorname{InjRad}_{X}}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
As the $\gamma$-dependency is now eliminated, the sum contributes the number of elements it sums over as a coefficient, which leads to the following equal expression:
$4\pi^{2}\,\frac{e^{8T}}{\operatorname{InjRad}_{X}^{2}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
Hence we get the following bound for Expression (6.10), as now we have a bound for the second factor:
$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\,\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz.$
By writing $z^{\prime}$ using polar coordinates around $z$, whence $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, we get the following equal expression:
$2\pi\,\|a\|_{\infty}^{2}\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho d\theta dz,$
which then equals
$4\pi^{2}\,\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho,$
as there was no $\theta$ dependency nor $z$ dependency.
The following lemma bounds a complicated integral in which $F_{t,t^{\prime},\rho}$ appears. The bounds for $F_{t,t^{\prime},\rho}$ are established in Section 7.
@@
Let $\beta$ be a positive constant that is smaller than $1$. Then:
$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}d\rho\leq\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
where
$F_{t,t^{\prime},\rho}\leq 32\pi\,\cdot\,\begin{cases}&\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad\text{ if }\rho<t-t^{\prime}\\ &\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},\qquad\text{ if }t-t^{\prime}<\rho<t\\ &\frac{1}{\sqrt{\sinh(\rho)}},\qquad\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
and where
$C_{6}=480\pi.$
The proof uses bounds for $F_{t,t^{\prime},\rho}$. Acquiring good enough bounds for the integral to converge requires very careful analysis which is done in Section 7. There we acquire three different bounds for $F_{t,t^{\prime},\rho}$ working in separate cases. Hence to bound the integral, we split it into three cases and use the bounds for $F_{t,t^{\prime},\rho}$. One of the cases is more technical as both the Beta function and the Gamma function appear in the analysis.
@@
We have that
$(1+\rho)\leq\frac{2}{\beta}\,e^{\frac{\beta}{2}-1}\,e^{-\frac{\beta}{2}\rho},$
which can be verified by studying the derivatives of the functions. Using this observation gives the following upper bound for the integral we are bounding:
$\frac{1}{\beta}\,\int_{0}^{2T}\sinh(\rho)\,e^{-\frac{\beta}{2}\rho}\,F_{t,t^{\prime},\rho}^{2}d\rho.$
By splitting the integral and using the bounds for $F_{t,t^{\prime},\rho}$, we have the following upper bound for the integral expression we set out to bound:
$\frac{32\pi}{\beta}\left(\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{\sqrt{\sinh(\rho)}\,e^{-\frac{\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\ +\ \int_{t-t^{\prime}}^{t}\frac{e^{-\frac{\beta}{2}\rho}}{(1-e^{t-t^{\prime}-\rho})^{\frac{1}{4}}}d\rho\ +\ \int_{t}^{t+t^{\prime}}e^{-\frac{\beta}{2}\rho}d\rho\right).$ (6.11)
Let us focus on the first term inside the parentheses. It is bounded by
$\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\int_{0}^{t-t^{\prime}}\frac{e^{\frac{1-\beta}{2}\rho}}{\sqrt{t-t^{\prime}-\rho}}d\rho\leq\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\int_{0}^{t-t^{\prime}}\frac{1}{\sqrt{t-t^{\prime}-\rho}}d\rho,$
where the inequality followed from using the upper boundary of the integration domain. The previous then equals after evaluating the integral:
$2\,\frac{e^{\frac{t^{\prime}-t}{2}}}{\sqrt{\sinh(t-t^{\prime})}}\,e^{\frac{1-\beta}{2}(t-t^{\prime})}\,\sqrt{t-t^{\prime}}\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
The inequality followed from the observation that a positive $x$ divided by $\sinh(x)$ is bounded from above by $1$.
Next, let us bound the second term inside the parentheses of Expression (6.11). We begin by writing it in a different form using a change of variables $u=t^{\prime}+\rho-t$:
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{t^{\prime}}e^{-\frac{\beta}{2}u}\,\left(1-\frac{1}{\cosh(u)}\right)^{-\frac{1}{4}}\,du.$
We do another change of variables $z=e^{-u}$ to get
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,\frac{dz}{z},$
where especially
$1-\frac{1}{\cosh(u)}=\frac{(1-z)^{2}}{1+z^{2}}.$
The previous integral expression equals
$e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{e^{-t^{\prime}}}^{1}z^{\frac{\beta}{2}-1}\,(1+z^{2})^{\frac{1}{4}}\,(1-z)^{-\frac{1}{2}}\,dz\leq 2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,\int_{0}^{1}z^{\frac{\beta}{2}-1}\,(1-z)^{-\frac{1}{2}}\,dz,$
where the inequality followed from that $(1+z^{2})^{\frac{1}{4}}$ is bounded by $2$ and the integration domain is enlargened. The previous expression then equals
$2\,e^{-\frac{\beta}{2}(t-t^{\prime})}\,B\left(\frac{\beta}{2},\frac{1}{2}\right),$ (6.12)
where $B$ is the Beta function which can be written using Gamma functions (the reference for the Beta and Gamma functions appearing here is [Olv10, Chp.6]):
$B\left(\frac{\beta}{2},\frac{1}{2}\right)=\frac{\Gamma\left(\frac{1}{2}\right)\,\Gamma\left(\frac{\beta}{2}\right)}{\Gamma\left(\frac{1}{2}+\frac{\beta}{2}\right)}.$
We have that $\Gamma\left(\frac{1}{2}\right)$ equals $\sqrt{\pi}$. We also have that
$\Gamma(x)=\frac{1}{x}\Gamma(x+1),$
and that
$\frac{\Gamma(x+1)}{\Gamma\left(x+\frac{1}{2}\right)}\leq(x+1)^{\frac{1}{2}},$
for positive $x$. Hence
$B\left(\frac{\beta}{2},\frac{1}{2}\right)\leq\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}.$
Hence Expression (6.12) is bounded by
$2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})}.$
Let us focus then on the third term inside the parentheses of Expression (6.11). The integral can be evaluated thus giving the following equal expression:
$\frac{2}{\beta}\left(e^{-\frac{\beta}{2}t}-e^{-\frac{\beta}{2}(t+t^{\prime})}\right),$
which has the following upper bound
$\frac{2\,e^{-\frac{\beta}{2}(t-t^{\prime})}}{\beta}.$
All three terms inside the parentheses of Expression (6.11) are bounded and they give the final upper bound for the original integral:
$\frac{32\,\pi}{\beta}\,e^{-\frac{\beta}{2}(t-t^{\prime})}\left(2\ +\ 2\sqrt{\pi}\,\frac{2}{\beta}\,\left(1+\frac{\beta}{2}\right)^{\frac{1}{2}}\ +\ \frac{2}{\beta}\right)\leq\frac{480\pi}{\beta^{2}}\,e^{-\frac{\beta}{2}(t-t^{\prime})},$
where we especially noted that $\beta<1$ for the inequality.

### v2

@@ 6. Geometric data
$\int_{\mathbb{H}}a(z^{\prime})^{*}\left(\int_{\mathbb{H}}K(x,z)L(x,z^{\prime})dx\right)\sum_{\gamma\in\Gamma}\left(\int_{\mathbb{H}}K(y,\gamma z)L(y,z^{\prime})dy\right)dz^{\prime}.$
The following lemma addresses integration over the part of the fundamental domain with large injectivity radius, yielding an upper bound consisting of two terms. The first term is an integral whose integrand exhibits exponential decay, while the second term is an integral over points with small injectivity radius. The latter term is bounded in Lemma 6.5.
@@
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$. Let $D(4T)$ denote the points with injectivity radius larger than $4T$. Let $a$ be a mean zero function on $X$. Then
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|$
= $=\ 30\|a\|_{2}^{2}\,\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}\,d\rho+\|a\|_{\infty}^{2}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz,$ (6.5)
where $\beta=\beta(\lambda_{1}(X))$, $\lambda_{1}(X)$ is the spectral gap of the Laplace–Beltrami operator $\Delta_{X}$, and
$\beta(x)=\begin{cases}1-\sqrt{1-4x},&\text{ if }x\leq\frac{1}{4},\\ 1,&\text{ if }x>\frac{1}{4}.\end{cases}$
The proof is based on noticing that only the identity element can have a non-zero contribution. After this, we separate the integration over the whole fundamental domain and the integration over points with small injectivity radius using the triangle inequality. The crucial part is bounding the first term using the exponential mixing theorem, Theorem 1.2.
@@
As $z\in D(4T)$, we have $d(z,\gamma z)>4T$ for any non-identity element $\gamma\in\Gamma$. Also, $z^{\prime}\in B(z,2T)$, so $d(\gamma z,z^{\prime})>2T$ for any non-identity element. Since $F_{t,t^{\prime},d(z,z^{\prime})}=0$ if $d(z,z^{\prime})>2T$, only the identity element contributes. Hence the left-hand side of Expression (6.5) equals
$\left|\int_{D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.6)
Adding and subtracting
$\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz$
and applying the triangle inequality gives the following upper bound for Expression (6.6):
$\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|\ +\ \left|\int_{D\setminus D(4T)}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|.$ (6.7)
We focus on the second term of the previous expression. By moving the absolute value inside, bounding $|a|$ by its supremum norm, and introducing a sum, we obtain an upper bound corresponding to the second term on the right-hand side of Expression (6.5).
Focus on the first term of Expression (6.7). Write $z^{\prime}$ in polar coordinates around $z$, so $z^{\prime}=\pi_{1}(\varphi_{\rho}(z,\theta))$, giving
$\left|\int_{D}a(z)\int_{\mathbb{S}^{1}}\int_{0}^{2T}\sinh(\rho)\,a(\varphi_{\rho}(z,\theta))^{*}\,F_{t,t^{\prime},\rho}^{2}\,d\rho\,d\theta\,dz\right|.$
Reordering the integrals using Fubini’s theorem gives
$\left|\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\int_{D}\int_{\mathbb{S}^{1}}a(z)\,a(\varphi_{\rho}(z,\theta))^{*}\,d\theta\,dz\,d\rho\right|.$
Pulling the absolute value inside the first integral, we have
@@
Applying the Exponential Mixing Theorem, Theorem 1.2, to the inner product yields the upper bound
$\|a\|_{2}^{2}\,\int_{0}^{2T}11\,\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,e^{\beta}(1+\rho)e^{-\beta\rho}\,d\rho\leq 30\|a\|_{2}^{2}\int_{0}^{2T}F_{t,t^{\prime},\rho}^{2}\,(1+\rho)\,e^{-\beta\rho}\,d\rho,$
where the latter inequality followed by using that $\beta\leq 1$ and $11e\leq 30$. This concludes the proof.
The following lemma bounds the integration over the part of the fundamental domain with small injectivity radius.
@@
Let $X=\Gamma\setminus\mathbb{H}$ be a compact connected hyperbolic surface with fundamental domain $D$, and let $D(4T)$ denote the points with injectivity radius larger than $4T$. Then
$\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\leq 4\pi^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}d\rho.$
The proof is based on the careful use of Hölder’s inequality. Of particular importance is understanding the number of group elements with non-zero contribution.
@@
$\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$ (6.8)
Only the terms $\gamma\in\Gamma$ with $d(\gamma z,z^{\prime})<2T$ for some $z^{\prime}\in B(z,2T)$ contribute. Denote this collection by $\Gamma(z)$. Then
$\#\Gamma(z)\leq\frac{|B(z,4T)|}{\operatorname{InjRad}_{X}}\leq 2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}.$
This is because $\gamma z$ must lie in $B(z,4T)$ for it to be possible that $d(\gamma z,z^{\prime})<2T$, as $z^{\prime}\in B(z,2T)$. Restricting the sum over $\Gamma(z)$ in Expression (6.8), and then applying Hölder’s inequality twice, first to the integral and then to the sum, results in the following upper bound:
$\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}\left(\int_{D\setminus D(4T)}\int_{B(z,2T)}\#\Gamma(z)\sum_{\gamma\in\Gamma(z)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}.$ (6.9)
Focus on the right factor. It admits the following upper bound after using Fubini’s theorem to reorder the sums and integrals, and then applying the bound on $\Gamma(z)$:
@@
$\left(\frac{2\pi\,e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(\gamma z,z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}.$
Changing variables via $z^{\prime}\mapsto\gamma z^{\prime}$ gives the following equal expression:
@@
$\left(\frac{2\pi\,e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(\gamma^{-1}z,2T)}F_{t,t^{\prime},d(\gamma z,\gamma z^{\prime})}^{2}dz^{\prime}dz\right)^{\frac{1}{2}}.$
Since $d(\gamma z,\gamma z^{\prime})=d(z,z^{\prime})$ and $F_{t,t^{\prime},\rho}=0$ for $\rho>2T$, the domain of integration in the inner integral can be restricted to $B(\gamma^{-1}z,2T)\cap B(z,2T)\subset B(z,2T)$. Hence, we obtain the following upper bound for the previous integral expression:
@@
$\left(\frac{2\pi\,e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\sum_{\gamma\in\Gamma(z)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,\gamma z^{\prime})}^{2}\,dz^{\prime}\,dz\right)^{\frac{1}{2}}$
@@
$\leq\ \left(\frac{4\pi^{2}\,e^{8T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,\gamma z^{\prime})}^{2}\,dz^{\prime}\,dz\right)^{\frac{1}{2}}.$
The inequality follows because the contents of the sum no longer depends on $\gamma$, so we can use the bound on $\Gamma(z)$.
Hence, we obtain the following upper bound for Expression (6.9):
$2\pi\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}^{2}\,dz^{\prime}\,dz=4\pi^{2}\,|D\setminus D(4T)|\,\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)\,F_{t,t^{\prime},\rho}^{2}\,d\rho,$
where the equality follows from writing $z^{\prime}$ in polar coordinates around $z$ and then integrating over $z$. This concludes the proof.
The following lemma bounds an integral in which $F_{t,t^{\prime},\rho}$ appears.
@@
Let $\beta$ be a positive constant smaller than $1$ and let $0<t^{\prime}<t$. Then
$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}d\rho\leq\frac{C_{6}}{\beta^{2}}\,e^{-\frac{\beta}{4}(t-t^{\prime})},$
when
$F_{t,t^{\prime},\rho}\leq 50\frac{1}{\sqrt{\sinh(\max(t-t^{\prime},\rho))}}\bm{1}_{\rho\in(0,t+t^{\prime})},$
where $C_{6}=10000$.
The proof is based on splitting the integral into two terms according to the upper bound of $F_{t,t^{\prime},\rho}$.
@@
We seek to bound
$\int_{0}^{2T}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}\,F_{t,t^{\prime},\rho}^{2}\,d\rho$
$\leq 50^{2}\left(\frac{1}{\sinh(t-t^{\prime})}\int_{0}^{t-t^{\prime}}\sinh(\rho)\,(1+\rho)\,e^{-\beta\rho}\,d\rho+\int_{t-t^{\prime}}^{t+t^{\prime}}(1+\rho)\,e^{-\beta\rho}\,d\rho\right),$
where the inequality follows from the bound on $F_{t,t^{\prime},\rho}$ and splitting the integration domain. Using
$(1+\rho)\leq\frac{2}{\beta}\,e^{\frac{\beta}{2}-1}\,e^{\frac{\beta}{2}\rho}\leq\frac{2}{\beta}e^{\tfrac{\beta}{2}\rho},$
which can be verified by analyzing derivatives, we obtain the following upper bound for the previous integral expression:
$\frac{2\cdot 50^{2}}{\beta}\left(\frac{1}{\sinh(t-t^{\prime})}\int_{0}^{t-t^{\prime}}\sinh(\rho)\,e^{-\frac{\beta}{2}\rho}d\rho+\int_{t-t^{\prime}}^{t+t^{\prime}}e^{-\frac{\beta}{2}\rho}d\rho\right).$ (6.10)
We have
$\sinh(\rho)\,e^{-\tfrac{\beta}{2}\rho}\leq\sinh(t-t^{\prime})\,e^{-\tfrac{\beta}{2}(t-t^{\prime})}$
in the first term of the previous integral expression. Applying this inequality to the first term of Expression (6.10) and then integrating both integrals yields the following upper bound for Expression (6.10):
$\frac{2\cdot 50^{2}}{\beta}\left((t-t^{\prime})\,e^{-\tfrac{\beta}{2}(t-t^{\prime})}+\frac{2}{\beta}\left(e^{-\tfrac{\beta}{2}(t-t^{\prime})}-e^{-\tfrac{\beta}{2}(t+t^{\prime})}\right)\right).$ (6.11)
Using that
$xe^{-bx}\leq\frac{2}{e}\frac{1}{b}e^{-bx/2}$
for all $x>0,b\in(0,1)$, Expression (6.11) is bounded by
$\frac{10000}{\beta^{2}}\,e^{-\frac{\beta}{4}(t-t^{\prime})}.$